OpenStudy (anonymous):

I really need help understanding this! 1. Which two points lie on a line with slope 2/3? (2, 0) and (3,0) (3, –4) and (1, 2) (0, 0) and (2, 3) (–3, 2) and (0, 4) 2. What is the slope of the line that passes through the given points? (6, –1) and (–3, –1) undefined 0 – 2/3 – 3/2 4. Write the equation in slope-intercept form. What are the slope and y-intercept? 3/4x + 4/5y = 4 y = 15/16x + 5; slope: 5; y-intcerpt: 15/16 y = 15/16x – 5; slope: 15/16 : y-intercept: –5 y = –15/16x – 5; slope: –15/16; y-intercept: –5 y = –15/16x + 5; slope: –15/16; y-intercept: 5

OpenStudy (anonymous):

1. Which two points lie on a line with slope 2/3? use: m=(y2-y1)/(x2-x1) and plug in m which is slope and see which of the choices satisfies the equation 2/3=(y2-y1)/(x2-x1) It is (–3, 2) and (0, 4) (2-4)/(-3-0)=-2/-3=2/3 :)

OpenStudy (anonymous):

It's still confusing o.O

OpenStudy (anonymous):

do you go to connections academy

OpenStudy (anonymous):

No but my schooling uses the same things as connections academy such as connexus and stuff like that :)

OpenStudy (anonymous):

i recogizne the test that why i would help but i need unit name

OpenStudy (anonymous):

Oh Functions, Equations, and Graphs : Linear functions and slope intercept.

OpenStudy (anonymous):

All you have to do for q 1 is use the slope formula which is. m=(y2-y1)/(x2-x1) where m is slope and it will be 2/3 y2 is 2nd y y1 is first y x2 is 2nd x x1 is 1st x and just start pluging in.

OpenStudy (anonymous):

ohhhhhhhh okay. Got it.

OpenStudy (anonymous):

ook :D 2. What is the slope of the line that passes through the given points? (6, –1) and (–3, –1) you use the same formula y2 is 2nd y y1 is first y x2 is 2nd x x1 is 1st x 

OpenStudy (anonymous):

4. Write the equation in slope-intercept form. What are the slope and y-intercept? 3/4x + 4/5y = 4 I think you skipped number 3. But the above is just solve for y and you effectively put it in y intercept form.

OpenStudy (anonymous):

Thank you so much!!! :) Uhm I skipped 3 because I already got that one :P But thank you!

OpenStudy (anonymous):

De nada :)

OpenStudy (anonymous):

Woaaaa wait I still kinda don't get 4 o.O

OpenStudy (anonymous):

simply solve for y 3/4x + 4/5y = 4 4/5y = 4-3/4x 4/5y = 5/4*(4-3/4x) y=5/4*(4-3/4x) y = -(15/16)*x+5 distriute 5/4 and rearrange

OpenStudy (anonymous):

Oh okayyy! I tend to think things are more difficult than they really are. But thanks again :)

OpenStudy (anonymous):

Just take it slow and read carefully :D All things are very simple that way.

OpenStudy (anonymous):

Yea, sure thing :)